# A neural-network based nonlinear non-intrusive reduced basis method with online adaptation for parametrized partial differential equations

### v1

@@ 5 Conclusions
In this work, we have proposed a novel NN-based nonlinear RB method with online adaptation for efficient numerical solution of parametrized PDEs. The method integrates the strengths of non-intrusive RB approaches and PINNs, enabling both fast offline training and accurate online refinement. Unlike many existing methods that rely on linear combinations of basis functions, our approach employs a nonlinear mapping that better captures complex solution manifolds. The online adaptation stage leverages lightweight PINN-style training to refine predictions in an unsupervised manner, ensuring robustness and improving accuracy, even when offline-trained NN models are insufficient. Through numerical experiments on both a parametrized Burgers’ equation and a multi-parameter lid-driven cavity problem, we have demonstrated that the proposed method achieves high accuracy with significantly reduced computational cost. In particular, the online stage has proved especially important in scenarios involving complex solution structures or limited snapshot data, where it has substantially enhanced the reliability and efficiency of predictions.
@@ Appendix A Implementation details
Code and data : https://github.com/hzjy2025/nonintrusive25. This code is based on and extends the given code by [8].
In both numerical examples, we use same the network architectures, which are detailed below.
The dimension reduction model $\Phi$ in (3.1) consists of one fully connected layer with Swish activation [25], followed by an output layer with no activation. This network has $\ell$ input neurons, $r$ neurons at the hidden layer and $r$ outputs. We fix $\ell=20$ in our experiments and study the performance of the model for different values of $r$.
The solution reconstruction model $U$ (see (3.2)) is a one hidden layer network with $r$ inputs, $5$ hidden neurons and one output. The activation used is Tanh [25]. We emphasize that this network is deliberately kept simple as it needs to be trained online.
The input dimension of the parameter estimation model $\Theta$ (see (3.3)) is equal to the dimension of the parameter domain; this model has four hidden layers with $100$ neurons each. The activation after each hidden layer is Swish. The number of outputs of $\Theta$ is the total number of network parameters in $U$. Specifically, $U$ is represented by two weight matrices and two bias vectors, so the total number of network parameters is $5r+11$. In the lid-driven cavity example where we want to predict the pressure and two components of the velocity field, we use three models with the same structure as the one discussed above. Each model corresponds to one output.
In both test examples, the networks are trained in the offline stage using the loss function (3.5) with $p=1$. In the first example, we set $\lambda_{0}=1$ and $\lambda_{1}=0.001$ in (3.5). In the second example where the PDE model is two-dimensional, we set $\lambda_{(0,0)}=1$ and $\lambda_{(1,0)}=\lambda_{(0,1)}=0.001$.
In the second example, the pressure is determined up to an additive constant; see, e.g. [18]. To account for this, we compute the integral of the pressure function over the domain using the Clenshaw–Curtis quadrature and shift the function so that it has zero mean. We do this for all computed pressure functions, whether obtained using the proposed method or a standard PDE solver, and then measure the error between the shifted functions.

### v2

@@ 5 Conclusions
In this work, we have introduced an NNO framework as a nonlinear generalization of RB methods for parametrized steady-state PDEs. The method replaces the standard linear combination of basis functions with a structured, nonlinear mapping realized by a pointwise Nemytskii operator acting on fixed feature functions. By integrating nonlinear dimension reduction in the offline stage with lightweight, physics-informed online adaptation, the NNO framework accurately captures complex solution manifolds that cannot be well approximated by linear subspaces. Our numerical experiments on the one-dimensional Burgers’ equation and a three-dimensional nonlinear heat conduction problem demonstrate that the proposed method consistently outperforms the linear RB approach. Specifically, the online adaptation stage proves highly effective, yielding substantial improvements in prediction accuracy, particularly in parameter regimes where training data are limited.
@@ Appendix A Implementation details
In both numerical examples, we use the same network architectures, which are detailed below. The dimension reduction model $\mathbf{g}$ in (3.1) consists of one fully connected layer with Swish activation [25], followed by an output layer with no activation. This network has $r^{\prime}$ input neurons, $r$ neurons at the hidden layer and $r$ outputs. We fix $r^{\prime}=20$ in our experiments and study the performance of the model for different values of $r$. The solution reconstruction model $h$ (see (3.3)) is a one hidden layer network with $r$ inputs, $5$ hidden neurons and one output. The activation used is Tanh [25]. We emphasize that this network is deliberately kept simple as it needs to be trained online. The input dimension of the hypernet model $\sigma$ (see (3.4)) is equal to the dimension of the parameter domain; this model has four hidden layers with $100$ neurons each. The activation after each hidden layer is Swish. The number of outputs of $\sigma$ is the total number of network parameters in $h$. Specifically, $h$ is represented by two weight matrices and two bias vectors, so the total number of network parameters is $5r+11$.